# Learning out-of-time-ordered correlators with classical kernel methods

### v1

@@ Abstract
Out-of-Time Ordered Correlators (OTOCs) are widely used to investigate information scrambling in quantum systems. However, directly computing OTOCs with classical computers is often impractical. This is due to the need to simulate the dynamics of quantum many-body systems, which entails exponentially-scaling computational costs with system size. Similarly, exact simulation of the dynamics with a quantum computer (QC) will generally require a fault-tolerant QC, which is currently beyond technological capabilities. Therefore, alternative approaches are needed for computing OTOCs and related quantities. In this study, we explore four parameterised sets of Hamiltonians describing quantum systems of interest in condensed matter physics. For each set, we investigate whether classical kernel methods can accurately learn the XZ-OTOC as well as a particular sum of OTOCs, as functions of the Hamiltonian parameters. We frame the problem as a regression task, generating labelled data via an efficient numerical algorithm that utilises matrix product operators to simulate quantum many-body systems, with up to 40 qubits. Using this data, we train a variety of standard kernel machines and observe that the best kernels consistently achieve a high coefficient of determination ($R^{2}$) on the testing sets, typically between 0.9 and 0.99, and almost always exceeding 0.8. This demonstrates that classical kernels supplied with a moderate amount of training data can be used to closely and efficiently approximate OTOCs and related quantities for a diverse range of quantum many-body systems.
@@ I Introduction
OTOCs have also been used in discussions of quantum chaos [24, 25, 26], including in the study of black holes [27], thermalisation [28, 29], and many-body scarred systems [30]. The early-time exponential growth of the OTOC was initially thought to be a diagnosis of chaos, and was even used to define a quantum analogue of the Lyapunov exponent [31] and the Lyapunov spectrum [32]. However, more recent findings have revealed that the exponential growth of OTOCs is necessary but not sufficient to diagnose chaos in quantum systems [33], suggesting that this connection requires further research. This, together with the significant role that the OTOC plays in studying quantum information scrambling and other intriguing applications [34, 35], provides compelling reasons to undertake detailed investigations of the OTOC.
@@
On the other hand, procedures making use of quantum devices to directly compute OTOCs have been proposed [45, 46, 47, 48, 49]. However many of these procedures require either specialised quantum systems to simulate specific Hamiltonians, or access to a universal fault-tolerant quantum computer (QC). In the latter case one would, in principle, be able to simulate arbitrary quantum many-body systems. However, the devices currently in operation are noisy intermediate-scale quantum (NISQ) devices [50], which cannot reliably operate at the large scales required in the aforementioned procedures. Accordingly, we need to explore new methods for accurately and efficiently determining OTOCs, and related quantities, in order to support continued research in this direction.
@@
In this work we investigate whether classical kernel methods (KMs) [51, 52, 53], powerful machine learning (ML) algorithms used for data analysis tasks, may help provide a solution to this problem. A previous work [54] applied restricted Boltzmann machines to learn early-time OTOCs, and gave one example of the procedure for the 2D transverse-field Ising model. Here we consider four parameterised sets of Hamiltonians, each containing a collection of 1D quantum systems which are studied in the condensed matter physics literature. For each set, we investigate whether classical KMs can be used to learn a specific OTOC, namely the XZ-OTOC, and a useful sum of OTOCs (defined in Section II.3) as functions of the Hamiltonian parameters. We formulate the problem as a regression task, generating labelled data with an efficient numerical algorithm based on matrix product operators [41], for quantum many-body system sizes up to 40 qubits. The input data is given by uniform random samples of the Hamiltonian parameters drawn from a subset of the parameter space. The labels are given by either the value of the associated XZ-OTOC, or the sum of OTOCs. We split the generated data into training and testing sets and apply regularised empirical risk minimisation (RERM) [52, 53] with KMs. Using a variety of standard kernels, we perform a 10-fold cross-validation on the training data to find good hyperparameter values, and apply the best models to the testing data to assess their suitability for the problem.
@@
The paper proceeds as follows. In Section II, we explain the necessary background material for discussing the ML tasks in detail, followed by a formal introduction to OTOCs and the associated sum of OTOCs on which we focus. In Section III, we describe the parameterised sets of Hamiltonians considered in this work and the associated ML tasks, together with the classical kernels which are applied to the ML problem. We then finish this section by discussing the methods used to generate the datasets. In Section IV, we report the numerical results, which include learning performance metrics for the best models making predictions on the training and testing sets for every problem instance. In Section V, we discuss the results and their significance. Finally, in Section VI, we conclude and provide suggestions for improving and extending our work in future research.
@@ II.1 Regularised empirical risk minimisation
where $L_{\mathcal{D}}:\textsc{Hyp}(\mathcal{X},\mathbb{R})\to\mathbb{R}$ is a loss function for $\mathcal{D}$, $\Omega:\textsc{Hyp}(\mathcal{X},\mathbb{R})\to\mathbb{R}$ is the regularisation term, and $\lambda>0$ is the regularisation strength. A common choice of the loss function for regression is given by the mean squared error (MSE), $L_{\mathcal{D}}(f)=\frac{1}{M}\sum_{i=1}^{M}\left(f(\mathbf{x}_{i})-y_{i}\right)^{2}$. Similarly, common choices of the regularisation term include the $l_{1}$-norm and the squared $l_{2}$-norm of the model parameters, called Lasso and Tikhonov regularisation, respectively.
@@ II.2 Kernel methods
When selecting an algorithm for a supervised ML task, several factors need to be considered. These include the dataset size and complexity, the type of problem (i.e., regression, classification, etc.), the efficiency of the training algorithm, and the availability of tools for implementation. KMs [52, 51, 53] are a collection of widely used ML algorithms aimed at capturing complex patterns in moderately sized datasets. The success of KMs stems largely from the use of kernel functions, which implicitly compute inner products between embeddings of input data in high-dimensional feature spaces. By utilising kernel functions, intricate non-linear structures in the original data can sometimes translate into standard linear functions in the feature space (see Figure 1). This simplification facilitates the learning process, making it considerably easier to identify the structures in the data, even with a small number of data points. Moreover, in many cases the training landscape for KMs is convex which enables the discovery of optimal model parameters via a deterministic procedure. This makes kernels a desirable candidate for a variety of supervised ML tasks.
@@
Performing RERM with KMs amounts to choosing an appropriate kernel $\mathcal{K}$ and using the corresponding RKHS $\mathcal{R}_{\mathcal{K}}$ as the hypothesis class, whose elements can be written as in (6). Optimisation over this space would appear to require searching for the best choice of infinitely many real coefficients $\{\alpha_{i}\}_{i\in\mathbb{N}}$. However, the representer theorem (Theorem 4.2 in [52] and Theorem 6.11 in [53]) reduces the problem to a finite-dimensional one. Suppose that we have a finite training dataset $\mathcal{D}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{M}\subseteq\mathcal{X}\times\mathbb{R}$ and a regularisation term $\Omega(f)=g\left(||f||_{\mathcal{R}_{\mathcal{K}}}\right)$, where $g:[0,\infty)\to\mathbb{R}$ is some strictly increasing function, and $||\cdot||_{\mathcal{R}_{\mathcal{K}}}$ is the norm induced by the inner product on $\mathcal{R}_{\mathcal{K}}$. Under these conditions, the representer theorem states that the functions in $\mathcal{R}_{\mathcal{K}}$ which minimise the regularised empirical risk functional can be expressed in the form
@@
Therefore, we can find the minimisers of the regularised empirical risk functional by searching over the finite-dimensional space of coefficients $\vec{\alpha}=(\alpha_{i})_{i=1}^{M}\in\mathbb{R}^{M}$. Further, if the regularised empirical risk functional is convex, then the optimal $\vec{\alpha}$ is unique and we can compute it deterministically. In this work, we use the MSE loss function and the regularisation term defined by $\Omega(f)=\frac{1}{M}||f||_{\mathcal{R}_{\mathcal{K}}}^{2}$ for all $f\in\mathcal{R}_{\mathcal{K}}$. In this case the regularised empirical risk functional is convex (see Appendix A.1). From this, it can be shown (see Appendix A.2) that the optimal $\vec{\alpha}\in\mathbb{R}^{M}$ is given by
@@
From (8), we see that determining the optimal $\vec{\alpha}\in\mathbb{R}^{M}$ requires the Moore-Penrose pseudoinverse of the $M\times M$ matrix $K^{2}+\lambda K$ to be known. The time complexity of calculating the pseudoinverse of a square matrix, which is the most computationally expensive part of this algorithm, is cubic in the number of rows or columns. This implies that the time complexity of training a model with this algorithm scales as $\mathcal{O}(M^{3})$ with the size of the training data set $M\in\mathbb{N}$. Once the optimal $\vec{\alpha}$ has been determined, (7) shows that the time-complexity of predicting the label for a new input scales as $\mathcal{O}(M)$. Accordingly, it is important for the size of the training dataset to be reasonably small, otherwise the time required for training (and making predictions) can become prohibitively large. Similarly, we need to be able to compute the kernel $\mathcal{K}$ efficiently, otherwise both calculating the kernel matrix entries and making predictions could be computationally expensive. The procedures for determining the optimal model (i.e. training) and making predictions for new input data are summarised in Algorithm 1 and Algorithm 2, respectively. These are the algorithms which we apply to the learning problems considered in this work.
Algorithm 1 Training a regression model with kernel methods
Inputs: Training dataset $\mathcal{D}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{M}$, kernel $\mathcal{K}$, regularisation strength $\lambda$
Output: Coefficients $\vec{\alpha}\in\mathbb{R}^{M}$ of the model in (7) that solves RERM for $\mathcal{D}$ over $\mathcal{R}_{\mathcal{K}}$
1:for $i=1,\ldots,M$ do
2:     for $j=1,\ldots,M$ do
3:         Set $K_{ij}:=\mathcal{K}(\mathbf{x}_{i},\mathbf{x}_{j})$
4:     end for
5:end for
6:Calculate $\vec{\alpha}$ according to (8)
7:return $\vec{\alpha}$
Algorithm 2 Making a prediction using a trained regression model with kernel methods
Inputs: Input data sample $x\in\mathcal{X}$, input training data $\{\mathbf{x}_{i}\}_{i=1}^{M}$, kernel $\mathcal{K}$, coefficients $\vec{\alpha}\in\mathbb{R}^{M}$ output by Algorithm 1
Output: Predicted label $y\in\mathbb{R}$ associated with the input data sample $x\in\mathcal{X}$
1:Set $y:=0$
2:for $i=1,\ldots,M$ do
3:     Set $y:=y+\alpha_{i}\mathcal{K}(x,\mathbf{x}_{i})$
4:end for
5:return $y$
@@ II.3 Information Spreading and Out-of-Time-Ordered Correlators
where $W_{k}(t)\equiv e^{iHt}W_{k}e^{-iHt}$ is the operator $W_{k}$ evolved in the Heisenberg picture, and $\langle\cdot\rangle_{\rho}$ denotes the expectation value measured from the state $\rho$. The right-hand side of (9) is a measure of the size of the commutator $[V_{j},W_{k}(t)]$ in the state $\rho$. If $\rho$ is the maximally mixed state $\mathbb{I}^{\otimes n}/2^{n}$, where $\mathbb{I}$ is the $2\times 2$ identity matrix, then the right-hand side is proportional to the square of the Frobenius norm of the operator $[V_{j},W_{k}(t)]$. Thus, we see that the size of this commutator bounds the sensitivity of the expectation value of $W_{k}(t)$ to the initial perturbation $V_{j}$. If the commutator is small, then Alice and Bob would need to repeat the procedure many times in order for Bob to reliably distinguish between the two cases, and hence determine which bit was sent.
@@
If the Hamiltonian $H$ only contains local (e.g. nearest-neighbour) interactions, then the number of sites upon which the nested commutators act non-trivially generally increases with the sum index $l$ in (10). The operator $W_{k}(t)$ is thus highly non-local in general. However, for small $|t|$, the terms in the sum with large support (i.e. large $l$) are insignificant and only become significant as $|t|$ increases. This gradual increase in the magnitude of the non-local terms describes how $W_{k}(t)$ spreads to become increasingly non-local over time. The extent to which $W_{k}(t)$ fails to commute with an operator $V_{j}$ then depends on the magnitude of the terms in (10) which do not simply contain the identity operator on qubit $j$ (i.e. those with support on qubit $j$).
These considerations suggest quantifying the spread of information in terms of the size of commutators between local operators in the Heisenberg picture. To this end, we define the OTOC as follows. Firstly, the squared-commutator OTOC [28, 57], denoted $C_{jk}(t)$, is the non-negative real number defined such that
@@
As mentioned above, the expression in (11) is proportional to the square of the Frobenius norm of the operator $[V_{j},W_{k}(t)]$, and thus captures the size of this commutator. Since $V_{j}$ and $W_{k}$ are both Hermitian and unitary, one can expand the commutators in (11) to write $C_{jk}(t)=1-F_{jk}(t)$, where $F_{jk}(t)$ is the real number given by
@@
The latter quantity, $F_{jk}(t)$, will be referred to as simply the OTOC [31, 14, 15]. Note that the OTOC and squared-commutator OTOC are sometimes defined via an expectation value measured from an arbitrary state. However, for the purposes of this paper we will take them to be defined by an expectation value measured from the maximally mixed state $\mathbb{I}^{\otimes n}/2^{n}$, which ensures that $F_{jk}(t)$ is real.
@@
The remainder of the system, $\mathcal{S}-q_{A}$, is prepared in the maximally mixed state, $\rho_{\mathcal{S}-q_{A}}=\mathbb{I}^{\otimes n-1}/2^{n-1}$. Since $\mathcal{S}$ is a closed system, the reference system $R_{1}$ remains maximally entangled with $\mathcal{S}$ as the system evolves in time. Initially, $R_{1}$ is only entangled with the subsystem $q_{A}$, but as $q_{A}$ interacts with the rest of the system, the entanglement with $R_{1}$ generally spreads into the non-local degrees of freedom of $\mathcal{S}$. One can track the spreading of entanglement by examining to what extent Bob is able to recover the entanglement with $R_{1}$ from another qubit, such as $q_{B}$, in $\mathcal{S}$.
In the Hayden-Preskill protocol, Bob is also given access to a reference system introduced to purify the subsystem $\mathcal{S}-q_{A}$ [7]. Since $\mathcal{S}-q_{A}$ begins in a maximally mixed state, it can be purified with another $(n-1)$-qubit reference system $R_{2}$, with which $\mathcal{S}-q_{A}$ forms $n-1$ Bell pairs,
$\left|\psi\right\rangle_{(\mathcal{S}-q_{A})\cup R_{2}}=\tfrac{1}{2^{(n-1)/2}}(\left|00\right\rangle+\left|11\right\rangle)^{\otimes(n-1)}.$ (14)
@@
Figure 3: Tensor network representation of the systems involved in the Alice-Bob quantum communication protocol. Initially, subsystems $R_{1}$ and $q_{A}$ form a Bell state, while subsystems $\mathcal{S}-q_{A}$ and $R_{2}$ form $n-1$ Bell states. The unitary operator $e^{-iHt}$ is then applied to the system $\mathcal{S}$. Bob’s aim is to detect and quantify the entanglement with $R_{1}$ that has spread from $q_{A}$ to $q_{B}$.
@@
Beyond simply determining how much information has reached $q_{B}$, in [60] an efficient decoding procedure was proposed for recovering the entanglement with $R_{1}$, by applying a unitary operation on $q_{B}\cup R_{2}$ to reconstruct a Bell state on $R_{1}\cup q_{B}$. The fidelity of the decoding protocol is closely related to the sum of OTOCs appearing in (17). Because of the importance of this sum in determining the spreading of quantum information through a quantum many-body system, in the following sections we will investigate how to compute its value in different systems using kernel machines.

### v2

@@ Abstract
Out-of-Time Ordered Correlators (OTOCs) are widely used to investigate information scrambling in quantum systems. However, directly computing OTOCs with classical computers is an expensive procedure. This is due to the need to classically simulate the dynamics of quantum many-body systems, which entails computational costs that scale rapidly with system size. Similarly, exact simulation of the dynamics with a quantum computer (QC) will either only be possible for short times with noisy intermediate-scale quantum (NISQ) devices, or will require a fault-tolerant QC which is currently beyond technological capabilities. This motivates a search for alternative approaches to determine OTOCs and related quantities. In this study, we explore four parameterised sets of Hamiltonians describing local one-dimensional quantum systems of interest in condensed matter physics. For each set, we investigate whether classical kernel methods (KMs) can accurately learn the XZ-OTOC and a particular sum of OTOCs, as functions of the Hamiltonian parameters. We frame the problem as a regression task, generating small batches of labelled data with classical tensor network methods for quantum many-body systems with up to 40 qubits. Using this data, we train a variety of standard kernel machines and observe that the Laplacian and radial basis function (RBF) kernels perform best, achieving a coefficient of determination ($R^{2}$) on the testing sets of at least 0.7167, with averages between 0.8112 and 0.9822 for the various sets of Hamiltonians, together with small root mean squared error and mean absolute error. Hence, after training, the models can replace further uses of tensor networks for calculating an OTOC function of a system within the parameterised sets. Accordingly, the proposed method can assist with extensive evaluations of an OTOC function.
@@ I Introduction
OTOCs have also been used in discussions of quantum chaos [24, 25, 26], including in the study of black holes [27], thermalisation [28, 29], and many-body scarred systems [30]. The early-time exponential growth of the OTOC was initially thought to be a diagnosis of chaos, and was even used to define a quantum analogue of the Lyapunov exponent [31] and the Lyapunov spectrum [32]. However, more recent findings have revealed that the exponential growth of OTOCs is necessary but not sufficient to diagnose chaos in quantum systems [33], suggesting that this connection requires further research. This, together with the significant role that the OTOC plays in studying quantum information scrambling, and other intriguing applications [34, 35, 36], motivates detailed investigations of the OTOC. In this work, we focus on methods to numerically approximate the OTOC, as a means to further explore its dynamics more efficiently.
@@
On the other hand, procedures making use of quantum devices to directly compute OTOCs have been proposed [46, 47, 48, 49, 50]. However many of these procedures require either specialised quantum systems to simulate specific Hamiltonians, or access to a universal fault-tolerant quantum computer (QC). In the latter case one would, in principle, be able to simulate arbitrary quantum many-body systems. However, the only devices currently in operation are noisy intermediate-scale quantum (NISQ) devices [51]. In the case of 1D quantum systems, we believe that NISQ technologies will require significant improvements before they can match the simulation capabilities of classical tensor network methods (see Section V). For this reason, we focus on OTOCs describing 1D systems and explore applications of classical machine learning (ML) methods for accurately and efficiently determining such OTOCs (and related quantities), in order to support continued research in this direction.
To this end, we investigate whether classical kernel methods (KMs) [52, 53, 54], powerful ML algorithms used for data analysis tasks, may help to reduce the cost of extensively evaluating OTOCs describing 1D systems. Specifically, we explore whether KMs can accurately approximate such OTOCs from a small amount of training data. While neural networks are typically more efficient to train with larger datasets, here we seek a method for making accurate inferences from a small amount of data, since producing the data can be expensive. The small size of the datasets then makes the use of KMs over neural networks favourable since, for a given choice of hyperparameters, KMs offer a deterministic training procedure, allowing us to avoid potential convergence issues that can arise with neural networks. Additionally, in contrast to the growing computational cost of tensor network methods with system size (see Section III.4), once we obtain a trained model using KMs, we can use it to predict new OTOC values in a time that scales linearly in just the number of training datapoints (i.e., independent of the system size).
@@
In this work, we consider four parameterised sets of Hamiltonians, each containing a collection of 1D quantum systems which are studied in the condensed matter physics literature. For each set, we investigate whether classical KMs can be used to learn a specific OTOC, namely the XZ-OTOC, and a useful sum of OTOCs (defined in Section II.3) as functions of the Hamiltonian parameters. We formulate the problem as a regression task, generating labelled data with classical tensor network methods based on matrix product operators (MPOs) [42], for quantum many-body system sizes up to 40 qubits. The input data is given by uniform random samples of the Hamiltonian parameters drawn from a subset of the parameter space (see Section III.1). And the labels are given by either the value of the associated XZ-OTOC, or the sum of OTOCs. We split the generated data into training and testing sets and apply regularised empirical risk minimisation (RERM) [53, 54] with KMs. Using a variety of standard kernels, we perform a 10-fold cross-validation (see Section 4.5 of [54]) on the training data to find good hyperparameter values. We then use these hyperparameter values to train models on all of the training data, and apply the models to the testing data to assess their suitability for the problem.
An application of this method could be in situations where one is interested in understanding the behaviour of the OTOC over large regions of a Hamiltonian parameter space. For example, this could be used to design a quantum system with a particular information scrambling behaviour. The idea is that instead of extensively evaluating the OTOC (using, e.g., MPO methods) for many different Hamiltonian parameters, one can perform a small number of evaluations of the OTOC and then use KMs to approximate the OTOC for the remaining parameter values of interest. The benefit of this approach is that after the models are trained, further evaluations of the model can be done more efficiently than continued evaluations using MPO methods. In Section V, we elaborate upon how this provides an overall reduction in computational costs.
A previous work [55] applied restricted Boltzmann machines to learn early-time OTOCs, demonstrating their approach with an example involving the 2D transverse-field Ising model. Given that our work focuses on 1D quantum systems, a direct comparison with [55] is challenging. Nonetheless we include this reference as it is the only closely related work of which we are aware.
@@
The paper proceeds as follows. In Section II, we explain the necessary background material for discussing the ML tasks in detail, followed by a formal introduction to OTOCs and the associated sum of OTOCs on which we focus. In Section III, we describe the parameterised sets of Hamiltonians considered in this work and the associated ML tasks, together with the classical kernels which are applied to the ML problem. We then finish this section by discussing the methods used to generate the datasets. In Section IV, we report the numerical results, which include learning performance metrics for the trained models making predictions on the training and testing sets for every problem instance. Finally, in Section V, we discuss the results, describe how an overall reduction in runtime can be achieved, and conclude in Section VI, providing suggestions for improving and extending our work in future research.
@@ II.1 Regularised empirical risk minimisation
where $L_{\mathcal{D}}:\textsc{Hyp}(\mathcal{X},\mathbb{R})\to\mathbb{R}$ is a loss function for $\mathcal{D}$, $\Omega:\textsc{Hyp}(\mathcal{X},\mathbb{R})\to\mathbb{R}$ is the regularisation term, and $\lambda>0$ is the regularisation strength. A common choice of the loss function for regression is given by the mean squared error (MSE), $L_{\mathcal{D}}(f)=\frac{1}{M}\sum_{i=1}^{M}\left(f(\mathbf{x}_{i})-y_{i}\right)^{2}$. Similarly, common choices of the regularisation term include the $l_{1}$-norm and the squared $l_{2}$-norm of the model parameters, called Lasso and ridge regression (see Section 11.3 of [54]), respectively.
@@ II.2 Kernel methods
When selecting an algorithm for a supervised ML task, several factors need to be considered. These include the dataset size and complexity, the type of problem (i.e., regression, classification, etc.), the efficiency of the training algorithm, and the availability of tools for implementation. KMs [53, 52, 54] are a collection of ML algorithms used to capture complex patterns in small to moderate sized datasets. The success of KMs stems largely from the use of kernel functions, which implicitly compute inner products between embeddings of input data in high-dimensional feature spaces. By utilising kernel functions, intricate non-linear structures in the original data can sometimes translate into standard linear functions in the feature space (see Figure 1).
Moreover, in many cases the training landscape for KMs is convex which, for a given choice of hyperparameters, enables the discovery of optimal model parameters via a deterministic procedure. However, the training procedure has a cubic runtime scaling in the number of training datapoints. In contrast, back-propagation in conventional neural networks has a runtime scaling linearly in the number of training datapoints. So for large datasets, neural networks may be preferable. However if one uses a small enough training dataset, such as those used in this work, then kernels offer benefits including a deterministic training procedure, which motivates our use of them here.
@@
Performing RERM with KMs amounts to choosing an appropriate kernel $\mathcal{K}$ and using the corresponding RKHS $\mathcal{R}_{\mathcal{K}}$ as the hypothesis class, whose elements can be written as in (6). Optimisation over this space would appear to require searching for the best choice of infinitely many real coefficients $\{\alpha_{i}\}_{i\in\mathbb{N}}$. However, the representer theorem (Theorem 4.2 in [53] and Theorem 6.11 in [54]) reduces the problem to a finite-dimensional one. Suppose that we have a finite training dataset $\mathcal{D}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{M}\subseteq\mathcal{X}\times\mathbb{R}$ and a regularisation term $\Omega(f)=g\left(\|f\|_{\mathcal{R}_{\mathcal{K}}}\right)$, where $g:[0,\infty)\to\mathbb{R}$ is some strictly increasing function, and $\|\cdot\|_{\mathcal{R}_{\mathcal{K}}}$ is the norm induced by the inner product on $\mathcal{R}_{\mathcal{K}}$. Under these conditions, the representer theorem states that the functions in $\mathcal{R}_{\mathcal{K}}$ which minimise the regularised empirical risk functional can be expressed in the form
@@
Therefore, we can find the minimisers of the regularised empirical risk functional by searching over the finite-dimensional space of coefficients $\vec{\alpha}=(\alpha_{i})_{i=1}^{M}\in\mathbb{R}^{M}$. Further, if the regularised empirical risk functional is convex, then the optimal $\vec{\alpha}$ is unique and we can compute it deterministically. In this work, we use the MSE loss function and the regularisation term defined by $\Omega(f)=\frac{1}{M}\|f\|_{\mathcal{R}_{\mathcal{K}}}^{2}$ for all $f\in\mathcal{R}_{\mathcal{K}}$. In this case the regularised empirical risk functional is convex (see Appendix A.1). From this, it can be shown (see Appendix A.2) that the optimal $\vec{\alpha}\in\mathbb{R}^{M}$ is given by
@@
From (8), we see that determining the optimal $\vec{\alpha}\in\mathbb{R}^{M}$ requires the Moore-Penrose pseudoinverse of the $M\times M$ matrix $K^{2}+\lambda K$ to be known. The time complexity of calculating the pseudoinverse of a square matrix, which is the most computationally expensive part of this algorithm, is cubic in the number of rows or columns. This implies that the time complexity of training a model with this algorithm scales as $\mathcal{O}(M^{3})$ with the size of the training data set $M\in\mathbb{N}$. Once the optimal $\vec{\alpha}$ has been determined, (7) shows that the time-complexity of predicting the label for a new input scales as $\mathcal{O}(M)$. Accordingly, it is important for the size of the training dataset to be reasonably small, otherwise the time required for training (and making predictions) can become prohibitively large. Similarly, we need to be able to compute the kernel $\mathcal{K}$ efficiently, otherwise both calculating the kernel matrix entries and making predictions could be computationally expensive.
@@ II.3 Information Spreading and Out-of-Time-Ordered Correlators
where $W_{k}(t)\equiv e^{iHt}W_{k}e^{-iHt}$ is the operator $W_{k}$ evolved in the Heisenberg picture under $H$ for $t$ units of time, and $\langle\cdot\rangle_{\rho}$ denotes the expectation value measured from the state $\rho$. The right-hand side of (9) is a measure of the size of the commutator $[V_{j},W_{k}(t)]$ in the state $\rho$. If $\rho$ is the maximally mixed state $\mathbb{I}^{\otimes n}/2^{n}$, where $\mathbb{I}$ is the $2\times 2$ identity matrix, then the right-hand side is proportional to the square of the Frobenius norm of the operator $[V_{j},W_{k}(t)]$. Thus, we see that the size of this commutator bounds the sensitivity of the expectation value of $W_{k}(t)$ to the initial perturbation $V_{j}$. If the commutator is small, then Alice and Bob would need to repeat the procedure many times in order for Bob to reliably distinguish between the two cases, and hence determine which bit was sent.
@@
If the Hamiltonian $H$ only contains local (e.g., nearest-neighbour) interactions, then the number of sites upon which the nested commutators act non-trivially generally increases with the sum index $l$ in (10). The operator $W_{k}(t)$ is thus highly non-local in general. However, for small $|t|$, the terms in the sum with large support (i.e., large $l$) are insignificant and only become significant as $|t|$ increases. This gradual increase in the magnitude of the non-local terms describes how $W_{k}(t)$ spreads to become increasingly non-local over time. The extent to which $W_{k}(t)$ fails to commute with an operator $V_{j}$ then depends on the magnitude of the terms in (10) which do not simply contain the identity operator on qubit $j$ (i.e., those with support on qubit $j$).
These considerations suggest quantifying the spread of information in terms of the size of commutators between local operators in the Heisenberg picture. To this end, we define the OTOC as follows. Firstly, the squared-commutator OTOC (SC-OTOC) [28, 58], denoted $C_{jk}(t)$, is the non-negative real number defined such that
@@
Since $V_{j}$ and $W_{k}$ are both Hermitian and unitary, one can expand the commutators in (11) to write $C_{jk}(t)=1-F_{jk}(t)$, where $F_{jk}(t)$ is the real number given by
@@
The latter quantity, $F_{jk}(t)$, will be referred to as simply the OTOC [31, 14, 15].
The OTOC and SC-OTOC are sometimes defined via an expectation value measured from an arbitrary state. However, throughout this paper we will focus on the case of the maximally mixed state, which ensures that $F_{jk}(t)$ is real. This choice is well-motivated in the literature on scrambling, since it provides a simple setting where the dynamics of information spreading can be studied in isolation from any pre-existing correlations in the state [15, 7]. The absence of correlations in $\rho=\mathbb{I}^{\otimes n}/2^{n}$ also greatly simplifies the numerical simulations described in Section III.4. Further, since the case of the maximally mixed state is related to the Frobenius norm of the commutator, it can be used to derive a bound which holds for any state $\rho$. Specifically, using Hölder’s inequality, we can show that
$\frac{1}{2}\big{\langle}[V_{j},W_{k}(t)]^{\dagger}[V_{j},W_{k}(t)]\big{\rangle}_{\rho}\leq 2^{n}\|\rho\|_{\infty}\,C_{jk}(t),$ (13)
where $C_{jk}(t)$ is the quantity from (11) defined with respect to the maximally mixed state, and $\|\rho\|_{\infty}$ is the operator norm of $\rho$ (i.e., the largest eigenvalue of $\rho$). Note that if $\rho$ is pure, then $\|\rho\|_{\infty}=1$. In general, it can also be bounded by the purity $\|\rho\|_{\infty}\leq\sqrt{\text{tr}(\rho^{2})}$. Thus we see, from (13), that the choice of the state $\mathbb{I}^{\otimes n}/2^{n}$ in (11) captures general properties of the information spreading, since then $C_{jk}(t)$ can be used to infer bounds for any other state.
Associated with the OTOC and the SC-OTOC are three distinct time regimes—short, intermediate, and long—each characterised by different time-dependent behaviours. In the short-time regime, the SC-OTOC grows, with the growth rate depending on the system’s characteristics. For chaotic systems, it is expected to grow exponentially. The intermediate-time regime marks a transition from growth to decay, occurring around the so-called scrambling time of the system. Finally, in the long-time regime, the SC-OTOC decays to a constant value, possibly with some superposed oscillations. All three time regimes have been studied for various reasons in the literature (see Section 2 of [26] for an overview and references). Accordingly, methods for calculating OTOCs in any of the three regimes may be of practical use, but we first need to know where the boundaries of each regime lies. For qubit models, such as those considered in this work, the scrambling time scales at most linearly in the size of the system, depending on the locality of the Hamiltonian (see the fourth paragraph of the Introduction in [31]). As a result, we know that the short-time regime and part of the intermediate-time regime (which contains the scrambling time) must occur before a time which is linear in the system size. We will make use of this in Section III.1.
@@
The remainder of the system, $\mathcal{S}\setminus q_{A}$, is prepared in the maximally mixed state, $\rho_{\mathcal{S}\setminus q_{A}}=\mathbb{I}^{\otimes n-1}/2^{n-1}$. Since $\mathcal{S}$ is a closed system, the reference system $R_{1}$ remains maximally entangled with $\mathcal{S}$ as the system evolves in time. Initially, $R_{1}$ is only entangled with the subsystem $q_{A}$, but as $q_{A}$ interacts with the rest of the system, the entanglement with $R_{1}$ generally spreads into the non-local degrees of freedom of $\mathcal{S}$. One can track the spreading of entanglement by examining to what extent Bob is able to recover the entanglement with $R_{1}$ from another qubit, such as $q_{B}$, in $\mathcal{S}$.
In the Hayden-Preskill protocol, Bob is also given access to a reference system introduced to purify the subsystem $\mathcal{S}\setminus q_{A}$ [7]. Since $\mathcal{S}\setminus q_{A}$ begins in a maximally mixed state, it can be purified with another $(n-1)$-qubit reference system $R_{2}$, with which $\mathcal{S}\setminus q_{A}$ forms $n-1$ Bell pairs,
$\left|\psi\right\rangle_{\left(\mathcal{S}\setminus q_{A}\right)\cup R_{2}}=\tfrac{1}{2^{(n-1)/2}}(\left|00\right\rangle+\left|11\right\rangle)^{\otimes(n-1)}.$ (15)
@@
Figure 3: Tensor network representation of the systems involved in the Alice-Bob quantum communication protocol. Initially, subsystems $R_{1}$ and $q_{A}$ form a Bell state, while subsystems $\mathcal{S}\setminus q_{A}$ and $R_{2}$ form $n-1$ Bell states. The unitary operator $e^{-iHt}$ is then applied to the system $\mathcal{S}$. Bob’s aim is to detect and quantify the entanglement with $R_{1}$ that has spread from $q_{A}$ to $q_{B}$.
@@
Beyond simply determining how much information has reached $q_{B}$, in [61] an efficient decoding procedure was proposed for recovering the entanglement with $R_{1}$, by applying a unitary operation on $q_{B}\cup R_{2}$ to reconstruct a Bell state on $R_{1}\cup q_{B}$. The fidelity of the decoding protocol is closely related to the sum of OTOCs appearing in (18). Because of the importance of this sum in determining the spreading of quantum information through a quantum many-body system, in the following sections we will investigate how to approximate its value in different systems using kernel machines.